Demand: N= -5p+770, Supply: N= 2.3p
P= the price in dollars. Find the price at which supply and demand are equal. At this price, how many televisions can be supplied and sold each week?
A.) $105.48; 252
B.) $105.48; 243
C.) $332.61; 765
D.) $285.19; 656

Determine if the given ordered triple is a solution of the system. (-5, 5, 2)
x-y+z=-2
x+y+z=2
x+y-z=12

Answer by Alan3354(69443)

Determine if the given ordered triple is a solution of the system. (-5, 5, 2)
x-y+z=-2
-5 -5 + 2 = -8 --> no
-----------
x+y+z=2
x+y-z=12